Table 1. Forecast Error Variance Decomposition in the Smets-Wouters (2007) Model

Each Shock’s Percent Contribution to Variance of

Horizon  Type of Shock  Output  Hours Inflation
4 Quarters  1-Monetary shock  9.1  11.1  4.6
2-Productivity shock  26.1  8.1  5.1
3-Investment shock  25.1  26.8  3.3
4-Risk premium shock  9.9  12.8  0.7
5-Exogenous spending shock  15.2  20.9  0.5
6-Price markup shock  26.1  8.1  5.1
7-Wage markup shock  6.5  12.9  42.9
All shocks  100.0  100.0  100.0
Shocks 4−7  39.6  53.9  86.9
10 Quarters  1-Monetary shock  5.9  8.0  5.3
2-Productivity shock  31.6  4.1  4.5
3-Investment shock  18.5  18.7  3.6
4-Risk premium shock  4.2  6.3  0.7
5-Exogenous spending shock  8.2  13.8  0.7
6-Price markup shock  11.1  11.4  34.0
7-Wage markup shock  20.5  37.7  51.1
All shocks  100.0  100.0  100.0
Shocks 4−7  44.0  69.2  86.5
1,000 Quarters  1-Monetary shock  2.3  3.4  4.6
2-Productivity shock  29.5  2.0  4.0
3-Investment shock  7.9  8.6  3.4
4-Risk premium shock  1.6  2.6  0.6
5-Exogenous spending shock  4.2  10.5  1.0
6-Price markup shock  6.4  6.2  28.6
7-Wage markup shock  48.2  66.7  57.8
All shocks  100.0  100.0  100.0
Shocks 4−7  60.3  86.0  88.0

Note: Output and hours are logged and detrended but not differenced.